# GroundCap: A Visually Grounded Image Captioning Dataset

### v1

@@ 1 Introduction
In contrast with works that focus solely on object grounding [20, 44, 31, 37], GroundCap introduces several key innovations. Our framework assigns unique identifiers to each object and unifies the grounding of objects, actions, and locations while maintaining consistent object identity throughout descriptions. This approach enables rich contextual descriptions that can reference multiple objects from different classes within a single text span and resolve complex references (e.g., linking pronouns like “They” to specific objects like “person-1” and “person-2”). Additionally, our system can ground spans of text to specific portions of background elements, moving beyond treating them as single entities.
Our main contributions can be summarized as follows: (1) a novel dataset of 52,016 images with grounded captions that maintain object identity across references; (2) a unified grounding framework that links objects, actions, and locations while handling background elements; (3) gMETEOR, a new metric that combines caption quality with grounding accuracy; (4) a strong baseline model based on fine-tuned Pixtral-12B; (5) comprehensive evaluation through automated metrics, human assessment, and ChatGPT-4o evaluation; The dataset and fine-tuned model are publicly available on Hugging Face.
The paper is organized as follows: Section 2 reviews related work on image captioning, object detection, and grounded captioning. Section 3 details the creation of GroundCap, addressing image selection, object detection, and grounded caption generation and human refinement. Section 4 introduces gMETEOR, a novel metric designed to assess the quality of grounded captions. Section 5 presents a baseline model for GroundCap. Section 6 outlines the results of human evaluation of GroundCap, baseline model and human-refined captions. Section 7 outlines current limitations. Section 8 presents the conclusions and future work.
@@ 2.1 Image Classification and Object Detection
Early object recognition datasets focused on image classification, requiring binary labels indicating object presence [23, 34]. The transition to more realistic scenarios came with datasets like Caltech 101 [11] and Caltech 256 [13]. A significant leap forward was made with ImageNet [9], containing millions of images across thousands of categories, enabling breakthroughs in deep learning for visual recognition [21]. Object detection datasets build upon classification by requiring localization of objects. Notable examples include PASCAL VOC [10] and MS COCO [26], which provide bounding box annotations for multiple object categories in natural images. The RefCOCO datasets [18, 43] further extended object localization by introducing natural language descriptions to guide object selection. Recent work [38] has focused on improving annotation quality of existing benchmarks.
@@ 2.2 Image Captioning
Image captioning datasets combine visual and textual information, requiring models to generate descriptive sentences for images. Flickr30k [42] and MS COCO Captions [26] are widely used benchmarks in this domain. These datasets, however, typically provide only general descriptions without explicit grounding to specific image regions or objects. Visual Genome [20] takes a step towards grounded captioning by providing several region-specific captions for each image, where each caption describes a particular area defined by a single bounding box. While this approach offers more localized descriptions, it differs from our task in GroundCap. Recent advances in Multimodal Large Language Models, such as Flamingo [2], LLaVA [27], Pixtral-12B [1] and Qwen-VL [4], have significantly improved image captioning capabilities. While some of these models, like Qwen-VL, provide visual grounding capabilities, they lack the ability to maintain consistent object identities across multiple references within a caption.
@@ 2.3 Action Recognition
Datasets for action recognition in images and videos, such as HICO-DET [7] and AVA [14], are relevant as they annotate human actions and interactions with objects. HICO-DET provides annotations for 600 human-object interaction categories in static images, while AVA offers dense annotations of 80 atomic visual actions in movie clips. IXMAS [40] uses multiple camera views to record actions, enabling viewpoint-independent action recognition. HMDB [22] provides a large collection of 51 human action categories with over 6,766 video clips extracted from various sources. The UCF dataset [28] specifically addresses the challenge of recognizing actions “in the wild”, featuring videos that capture real-world scenarios. ActionHub [45] provides 3.6 million action video descriptions spanning 1,211 common actions, designed to enrich the semantic understanding required for zero-shot action recognition.
@@ 3.3 Grounded Caption Generation
Quality control is implemented through an evaluation system based on precision ($P$), recall ($R$), and F1 score metrics. These values follow standard information retrieval definitions [32]: $P=\text{TP}/(\text{TP}+\text{FP})$; $R=\text{TP}/(\text{TP}+\text{FN})$; and F1 score as the harmonic mean of precision and recall. TP (true positives) represents correctly grounded objects, FP (false positives) are objects mentioned in the caption but not present in detections, and FN (false negatives) are detected objects omitted from the caption.
@@ 6.2 Evaluation Criteria
Using a 5-point Likert scale (1: Poor, 5: Excellent), evaluators assessed each caption on five criteria. Object Grounding Precision evaluates how accurately the caption uses grounding tags (<gdo>, <gda>, <gdl>), including correct tag type selection, verification of referenced objects, and semantic correctness of tagged text spans. Grounding Recall measures whether all detected objects are appropriately referenced in the caption. Description Accuracy focuses on the correctness of described actions, relationships, and scene details, ensuring the caption’s interpretation matches the visual content. Language Quality considers grammatical correctness, natural flow, and coherence across sentences. Finally, Overall Quality provides a holistic assessment combining grounding accuracy, description completeness, and linguistic quality.
@@ 6.5 Inter-annotator Agreement
To assess the reliability of human evaluations, we measured agreement between annotators using Krippendorff’s alpha coefficient [19]. Table 4 shows the agreement levels across different caption types and evaluation criteria. Annotations of Pixtral-generated captions showed the strongest agreement ($\alpha>0.6$ for all criteria), with particularly high consensus on object identification (0.84), description quality (0.86), and overall assessment (0.82). This suggests that Pixtral’s outputs have consistent, well-defined characteristics that evaluators could assess reliably.
@@ 8 Conclusions and Future Work
We introduced GroundCap, a novel dataset for grounded captioning that provides detailed descriptions of visual scenes grounded on detected objects, actions, and locations. The proposed new metric, gMETEOR, combines language quality with grounding accuracy, providing improved quality assessment. We presented a baseline model for grounded captioning on GroundCap and conducted a human evaluation study to assess the quality of our automatically generated, baseline model, and human-refined captions. Our baseline model achieves strong grounding recall (0.96) while maintaining good language quality (METEOR score of 0.23). Human evaluation demonstrates the effectiveness of our approach, with human-refined captions receiving high ratings across all criteria (4.34/5.0 overall quality). Manual inspection revealed that the lowest-performing captions are often frames where few or no objects were detected or the image contents were unclear (e.g. in blurry frames). Finally, we found that ChatGPT-4o evaluations strongly correlate with human judgments (0.78 Pearson correlation for overall quality), suggesting its potential as a scalable evaluation tool, especially given the poor correlation between traditional metrics and human assessment. These results establish GroundCap as a valuable resource for advancing research in visually grounded language generation.
Future work will explore the extension of GroundCap to include diverse visual content from existing image captioning datasets like MS COCO [26] or Flickr30k [42]. We plan to expand our ID-based grounding system to the captioning of multiple sequential frames.

### v2

@@ 1 Introduction
In contrast with works that focus solely on object grounding [3, 4], GroundCap introduces several key innovations. Our framework assigns unique identifiers to each object and unifies the grounding of objects, actions, and locations while maintaining consistent object identity throughout descriptions. This approach enables rich contextual descriptions that can reference multiple objects from different classes within a single text span and resolve complex references (e.g., linking pronouns like “They” to specific objects like “person-1” and “person-2”). Our system can ground spans of text to specific portions of background elements, moving beyond treating them as single entities.
Our main contributions can be summarized as follows: (1) a novel dataset of 52,016 images with grounded captions that maintain object identity across references; (2) a unified grounding framework that links objects, actions, and locations while handling background elements; (3) gMETEOR, a new metric that combines caption quality with grounding accuracy; (4) a strong baseline model based on fine-tuned Pixtral-12B; (5) comprehensive evaluation through automated metrics, human assessment, and ChatGPT-4o evaluation; The dataset and fine-tuned model are available on Hugging Face.
The paper is organized as follows: Section 2 reviews related work on image captioning, object detection, and grounded captioning. Section 3 details the creation of GroundCap. Section 4 introduces gMETEOR, a novel metric designed to assess the quality of grounded captions. Section 5 presents a baseline model for GroundCap. Section 6 outlines the results of human evaluation of GroundCap, baseline model and human-refined captions. Section 7 outlines current limitations. Section 8 presents the conclusions and future work.
@@ 2.1 Image Classification and Object Detection
The development of object recognition began with datasets like Caltech 101 [5], which established foundations for categorizing visual objects. A significant leap forward was made with ImageNet [6], containing millions of images across thousands of categories, enabling breakthroughs in deep learning for visual recognition [7]. Object detection datasets build upon classification by requiring localization of objects. Notable examples include PASCAL VOC [8] and MS COCO [1], which provide bounding box annotations for multiple object categories in natural images. The RefCOCO datasets [9, 10] further extended object localization by introducing natural language descriptions to guide object selection. Recent work [11] has focused on improving annotation quality of existing benchmarks.
@@ 2.2 Image Captioning
Image captioning datasets combine visual and textual information, requiring models to generate descriptive sentences for images. Flickr30k [12] and MS COCO Captions [1] are widely used benchmarks in this domain. These datasets, however, typically provide only general descriptions without explicit grounding to specific image regions or objects. Visual Genome [13] takes a step towards grounded captioning by providing several region-specific captions for each image, where each caption describes a particular area defined by a single bounding box. While this approach offers more localized descriptions, it differs from our task in GroundCap. Recent advances in Multimodal Large Language Models, such as Flamingo [14], LLaVA [15] and Pixtral-12B [16] have significantly improved image captioning capabilities. While some of these models, like Qwen-VL, provide visual grounding capabilities, they lack the ability to maintain consistent object identities across multiple references within a caption.
@@ 2.3 Action Recognition
Datasets for action recognition in images and videos, such as HICO-DET [17] and AVA [18], are relevant as they annotate human actions and interactions with objects. HICO-DET provides annotations for 600 human-object interaction categories in static images, while AVA offers dense annotations of 80 atomic visual actions in movie clips. IXMAS [19] uses multiple camera views to record actions, enabling viewpoint-independent action recognition. HMDB [20] provides a large collection of 51 human action categories with over 6,766 video clips extracted from various sources. The UCF dataset [21] specifically addresses the challenge of recognizing actions “in the wild”, featuring videos that capture real-world scenarios.
@@ 3.3 Grounded Caption Generation
Quality control is implemented through an evaluation system based on precision ($P$), recall ($R$), and F1 score metrics. These values follow standard information retrieval definitions: $P=\text{TP}/(\text{TP}+\text{FP})$; $R=\text{TP}/(\text{TP}+\text{FN})$; and F1 score as the harmonic mean of precision and recall. TP (true positives) represents correctly grounded objects, FP (false positives) are objects mentioned in the caption but not present in detections, and FN (false negatives) are detected objects omitted from the caption.
@@ 6.2 Evaluation Criteria
Using a 5-point Likert scale (1: Poor, 5: Excellent), evaluators assessed each caption on five criteria. Object Grounding Precision evaluates how accurately the caption uses grounding tags (<gdo>, <gda>, <gdl>), including correct tag type selection, verification of referenced objects, and semantic correctness of tagged text spans. Grounding Recall measures whether all detected objects are appropriately referenced in the caption. Description Accuracy focuses on the correctness of described actions, relationships, and scene details, ensuring the caption matches the visual content. Language Quality considers grammatical correctness, natural flow, and coherence across sentences. Finally, Overall Quality provides a holistic assessment combining grounding accuracy, description completeness, and linguistic quality.
@@ 6.5 Inter-annotator Agreement
To assess the reliability of human evaluations, we measured agreement between annotators using Krippendorff’s alpha coefficient. Table 4 shows the agreement levels across different caption types and evaluation criteria. Annotations of Pixtral-generated captions showed the strongest agreement ($\alpha>0.6$ for all criteria), with particularly high consensus on object identification (0.84), description quality (0.86), and overall assessment (0.82). This suggests that Pixtral’s outputs have consistent, well-defined characteristics that evaluators could assess reliably.
@@ 8 Conclusions and Future Work
This paper introduced GroundCap, a novel dataset for grounded captioning that provides detailed descriptions of visual scenes grounded on detected objects, actions, and locations using an unified grounding framework that maintains object identity across multiple references. The proposed gMETEOR which combines language quality with grounding accuracy, providing improved quality assessment. Additionally a baseline model for grounded captioning on GroundCap and conducted a human evaluation study to assess the quality of our automatically generated, baseline model, and human-refined captions. Our baseline model achieves strong grounding recall (0.96) while maintaining good language quality (METEOR score of 0.23). Human evaluation demonstrates the effectiveness of our approach, with human-refined captions receiving high ratings across all criteria (4.34/5.0 overall quality). Manual inspection revealed that the lowest-performing captions are often frames where few or no objects were detected or the image contents were unclear (e.g. in blurry frames). We found that ChatGPT-4o evaluations strongly correlate with human judgments (0.78 Pearson correlation for overall quality), suggesting its potential as a scalable evaluation tool, especially given the poor correlation between traditional metrics and human assessment. These results establish GroundCap as a valuable resource for advancing research in visually grounded language generation.
There are several limitations of our work. First, GroundCap is derived from MovieNet and consequently inherits biases toward cinematic content (scenes with professional lighting, camera angles, and compositions) that may not reflect the diversity of real-world images. Second, our object detection system, while effective, occasionally produces false positives or misses objects, which propagates errors to the caption generation process. Another limitation is that our current approach does not fully address the temporal aspect of visual understanding as the captions describe static frames rather than actions unfolding over time. This represents an important area for future development, particularly for video understanding applications.
Future work could explore the extension of GroundCap content from existing image captioning datasets like MS COCO [1] or Flickr30k [12], addressing potential biases toward cinematic content. Another direction for future research is extending the ID-based grounding system to handle multiple sequential frames. This would enable the development of temporally consistent grounded video descriptions that track objects and their interactions over time. The development of end-to-end models that can simultaneously detect objects, assign consistent IDs, and generate grounded captions without relying on separate object detection systems could also be explored.